### v2

# Generative Logic: A New Computer Architecture for Deterministic Reasoning and Knowledge Generation

Abstract
We present Generative Logic (GL), a deterministic architecture that starts from user-supplied axiomatic definitions (and, optionally, a list of simple facts for counterexample (CE) construction), written in a minimalist Mathematical Programming Language (MPL), and systematically explores their deductive neighborhood. Definitions are compiled into a distributed grid of simple Logic Blocks (LBs) that exchange messages; whenever the premises of an inference rule unify, a new fact is emitted with full provenance to its sources, yielding replayable, auditable proof graphs. A prototype software implementation instantiates the workflow on first-order Peano arithmetic. Starting only from the Peano axioms, GL enumerates conjectures, applies normalization, type, and CE filter, and automatically reconstructs machine-checkable proofs of foundational arithmetic laws, including associativity and commutativity of addition, associativity and commutativity of multiplication, and distributivity. On commodity hardware, the prover phase requires approximately 7 seconds; a complete run finishes in about 5 minutes. Generated proofs export to navigable HTML so that every inference step can be inspected independently. We outline a hardware-software co-design path toward massively parallel realizations and describe prospective integration with probabilistic models (e.g., large language models) for auto-formalization and conjecture seeding. The Python, C++, and MPL code to reproduce the Peano experiments, along with the full proof graphs in HTML as well as machine-readable text format, are available in the project’s GitHub repository at github.com/Generative-Logic/GL commit 56c9233 and are permanently archived at doi:10.5281/zenodo.17206386.
1 Introduction
The pursuit of automated mathematical reasoning currently stands at a crossroads, defined by two distinct paradigms. On one side, Large Language Models (LLMs) have demonstrated a remarkable capacity for solving routine problems and reformulating existing knowledge, yet their probabilistic nature limits their reliability in discovering novel, non-trivial proofs [11]. On the other side, interactive proof assistants such as Lean and Coq offer formal guarantees but remain fundamentally manual endeavors, requiring extensive human expertise and guidance [9]. This leaves the vast landscape of deep, creative mathematical exploration largely inaccessible to full automation.
This paper introduces Generative Logic (GL), a novel computer architecture conceived to bridge this gap. GL is designed to automate deep, definition-based mathematical reasoning by shifting the paradigm from human-assisted, single-theorem proving to the fully automatic generation and verification of entire families of theorems. At its core, the system takes a set of formal axioms and algorithmically explores the complete deductive space they define. This approach enables GL to systematically generate conjecture, produce machine-verifiable proofs, and iterate through complex logical chains at computational speeds.
The implications of such a system are significant. By accelerating the cycle of discovery and verification, GL has the potential to act as a catalyst for mathematical research and to shorten the path from abstract breakthroughs to downstream innovations in physics, engineering, and computer science. Furthermore, we envision GL serving as a deterministic reasoning core that can be integrated into probabilistic Artificial Intelligence (AI) workflows, providing a source of verifiable truth to enhance the capabilities of next-generation AI systems – a vision shared by recent work calling for the fusion of LLMs with formal methods to build more trustworthy AI [12]. In essence, just as the electronic calculator democratized arithmetic, GL aims to democratize high-level mathematical reasoning.
This paper presents the foundational principles, architecture, and early results of GL. We detail the system’s massively parallel, hardware-aware proof engine, its custom Mathematical Programming Language (MPL) for formalizing definitions, and its unique execution model. A case study in Peano arithmetic demonstrates the system’s ability to autonomously derive and prove foundational theorems, validating the architecture and its approach to automated discovery.
2 The GL Workflow
The GL system transforms a set of formal definitions into a web of proven theorems through a deterministic, multi-stage pipeline. This process is designed to be exhaustive, shifting the primary human contribution to the initial formulation of clear and consistent definitions. The workflow, outlined below, mirrors a software development cycle where errors in definitions can lead to "compilation" failures or run-time contradictions, which are themselves valuable for debugging the axiomatic foundation. Optionally, simple facts expressed in MPL (e.g., small addition/multiplication tables) can be supplied to filter out wrong conjectures and to reduce resource consumption.
  1. 1.
    Definition Input. The process begins with a user, or an LLM assistant, providing a set of foundational axioms and definitions. These are formalized using MPL, a domain-specific language designed to express second-order predicate logic in a machine-readable format. The integrity of the entire reasoning process depends on the quality of this initial input.
  2. 2.
    Conjecture Generation. Using the provided definitions, GL automatically and combinatorially generates a large set of conjectures. To manage the combinatorial explosion inherent in this step, GL employs a "weaving" process on a regularized theorem structure—typically of the form $A\implies(B\implies(C\implies D))$—and applies aggressive normalization and filtering techniques to discard redundant or malformed conjectures. This phase is purely syntactic and does not attempt to evaluate the semantic value of the conjectures.
  3. 3.
    Counterexample (CE) Filtering. A deterministic pre-proof triage stage that, for each conjecture, reuses GL’s prover architecture to test consistency against small arithmetic tables. Facts are woven with a peek-and-prune strategy to generate new facts based on the conjecture under test; any contradiction triggers immediate rejection.
  4. 4.
    Compilation and Disintegration. The generated conjectures are then compiled and distributed across the grid of Logic Blocks (LB). Each Logical Entity (LE) within a conjecture is assigned to a specific block, forming a distributed processing chain. Within each block, the relevant definitions are disintegrated into their constituent logical parts, which serve as the initial fuel for the proof search.
  5. 5.
    Iterative Proof Execution. With the system primed, GL begins a "flood" of asynchronous iterations to find proofs for all conjectures simultaneously. This core execution process is fundamentally symbolic and hash-table based. For a given implication, the premise acts as a hash request (the key) and the conclusion as the expected response (the value). During each computation cycle, a LB performs a burst of hash requests, combining its local expressions with incoming mail to probe the vast space of logical consequences.
    This process is deeply distributed; each LB operates independently on its local memory, and communication with other blocks only occurs between cycles. This design allows for massive parallelization across thousands or even millions of cores, turning a computationally intractable search into a manageable, parallel process.
  6. 6.
    Verifiable Output. Once the execution run is complete, GL generates a set of HTML files that serve as the final, verifiable output for visualization and debugging. These files contain a human-readable list of all proven theorems. Crucially, each step in every proof is hyperlinked to its justification—either an initial definition or a previously proven theorem—creating a fully auditable and transparent proof graph that can be independently verified.
3 System Architecture
The GL workflow is supported by a unique hardware-aware architecture designed for massive parallelism and deterministic symbolic manipulation. The system is not a monolithic processor but a distributed network of simple, independent processing elements that collectively perform complex reasoning tasks. This design can be realized physically as an Application-Specific Integrated Circuit (ASIC) or implemented in software on cloud infrastructure or multi-core servers.
3.1 Architecture Overview
The core of GL is a symbolic reasoning engine. Unlike traditional theorem provers that rely on heuristic search algorithms, GL treats logical inference as a memory-access problem. In this paradigm, an implication is modeled as a key-value pair in a distributed hash table. For an implication such as $(A\land B)\implies C$:
  • The premise, $(A\land B)$, serves as the hash key.
  • The conclusion, $C$, serves as the hash value.
Reasoning is thus transformed into a process of formulating hash requests from known logical expressions and, upon a successful lookup, receiving a new, valid expression as the response. Both the keys and values are represented as plain text strings, underscoring the symbolic nature of the architecture.
This engine is realized as a deeply distributed system of independent nodes, or LBs, which are connected through configurable, directed paths. Each LB possesses its own local memory and executes its core operation—a burst of hash requests—asynchronously. Newly generated expressions are passed to other nodes according to a pre-compiled connection scheme. This block-based architecture, where each node operates independently, is the key to the system’s scalability. It allows the computational load to be parallelized across thousands or even millions of cores, turning a search problem that would be intractable on a single processor into a manageable, massively parallel task.
Consider the following example. Full Implication:
$(n>m\land m>2)\implies(n>2)$
Premise:
$n>m\land m>2$
Conclusion:
$n>2$
Figure 1: An example for connection of two LBs.
Figure 1 visualizes in a simple way how two LBs can be connected.
3.2 MPL
The GL system requires a formal, machine-readable way to express mathematical definitions and axioms. To this end, we developed MPL, a language based on a constrained subset of second-order predicate logic. MPL is designed to be expressive enough to capture complex mathematical structures while remaining sufficiently simple for algorithmic analysis and compilation. This corresponds to the well-established logical framework of higher-order logic, as detailed in [1].
3.2.1 LEs and Ports
The fundamental unit in MPL is the LE, which represents a predicate that evaluates to a Boolean value. An LE can be conceptualized as a template function in a language like C++. Each LE has a set of "ports" through which it receives arguments. For example, the LE representing the statement $x>2$ has three ports: one for $x$, one for the constant $2$, and one for the operator >>.
Each port is defined by a definition set, which functions as a type system. For the LE $x>2$, the ports for $x$ and $2$ might have the definition set $\mathbb{N}$ (the set of natural numbers), while the port for >> would have the definition set $\mathcal{P}(\mathbb{N}\times\mathbb{N})$ (the set of binary relations on $\mathbb{N}$). This typed approach allows for strong compile-time checks on the validity of logical constructions.
3.2.2 Core Operators
New LEs are constructed from existing ones using a minimal set of three core operators, which are sufficient to express all of standard logic.
Negation (!)
The ! operator takes a single LE and inverts its Boolean output. For example, !(in[x,A]) corresponds to $x\notin A$. The ports and their definition sets remain unchanged.
Conjunction (&)
The (& A B) operator combines two LEs, $A$ and $B$, into a single LE that evaluates to true if and only if both $A$ and $B$ are true. Ports from $A$ and $B$ can be merged by assigning them the same reference name. For instance, in (& (in2[n,m,>]) (in2[m,q,<])), the port ‘m‘ is shared, effectively unifying the corresponding definition sets and creating a dependency between the two original LEs.
Implication (>>)
The (>[vars](A)(B)) operator is the most complex. If the variable list ‘[vars]‘ is empty, it represents logical implication, equivalent to $\neg A\lor B$. If the list is non-empty, such as in (>[n](A)(B)), it represents the universal quantifier "for all". This operator binds the variable ‘n‘ within the scope of the expression, asserting that for all possible values of ‘n‘ (as defined by its port’s definition set), the implication $\neg A(n)\lor B(n)$ holds.
The existential quantifier, $\exists$, is not needed as a primitive operator, as it can be constructed from negation and the universal quantifier: $\exists x:P(x)\equiv\neg(\forall x:\neg P(x))$.
3.2.3 Type System and Recursion Constraint
The MPL compiler enforces a strict type system when merging ports. A critical rule is the prevention of infinite recursive definitions. A definition set for a port cannot be unified with a set that contains it. For example, a port of type $X$ cannot be merged with a port of type $\mathcal{P}(X\times Y)$, as this would imply that the set $X$ contains a structure defined by itself. This restriction, analogous to type safety in programming languages, prevents paradoxes and ensures that all definitions are well-founded, though it limits the direct expression of concepts such as ordinal numbers.
3.2.4 Unique Naming of Reusable LEs
Complex, reusable concepts can be defined once and given a name. The listing below shows how the mathematical definition of a function $f:X\to Y$ is expressed in MPL. This LE, named ‘fXY‘, encapsulates the properties that for every element $x$ in the domain $X$, there exists a unique element $y$ in the codomain $Y$ with $f(x)=y$ .
1(fXY[f,X,Y]) :=
2(&
3 (>[x,y]
4 (in2[x,y,f])
5 (&
6 (in[x,X])
7 (in[y,Y])
8 )
9 )
10 (&
11 (>[x]
12 (in[x,X])
13 ! (>[y]
14 (in[y,Y])
15 ! (in2[x,y,f])
16 )
17 )
18 (>[x]
19 (in[x,X])
20 (>[y1,y2]
21 (&
22 (in2[x,y1,f])
23 (in2[x,y2,f])
24 )
25 (=[y1,y2])
26 )
27 )
28 )
29)
Listing 1: The definition of a function $f:X\to Y$ in MPL.
Another example of a foundational definition provided to the system, the Peano axioms for natural numbers, is shown below.
1(NaturalNumbers[N,i0,i1,s,+,*]) :=
2(&
3 (&
4 (in[i0,N])
5 (&
6 (in2[i0,i1,s])
7 (&
8 (fXY[s,N,N])
9 (&
10 (>[n]
11 (in[n,N])
12 !(in2[n,i0,s])
13 )
14 (>[m]
15 (in[m,N])
16 (>[n1,n2]
17 (&
18 (in2[n1,m,s])
19 (in2[n2,m,s])
20 )
21 (=[n1,n2])
22 )
23 )
24 )
25 )
26 )
27 )
28 (&
29 (&
30 (fXYZ[+,N,N,N])
31 (&
32 (>[a]
33 (in[a,N])
34 (>[b]
35 (in3[a,i0,b,+])
36 (=[a,b])
37 )
38 )
39 (>[b]
40 (in[b,N])
41 (>[a,c,d]
42 (&
43 (in2[b,c,s])
44 (in3[a,b,d,+])
45 )
46 (&
47 (>[e]
48 (in3[a,c,e,+])
49 (in2[d,e,s])
50 )
51 (>[e]
52 (in2[d,e,s])
53 (in3[a,c,e,+])
54 )
55 )
56 )
57 )
58 )
59 )
60 (&
61 (fXYZ[*,N,N,N])
62 (&
63 (>[a]
64 (in[a,N])
65 (>[b]
66 (in3[a,i0,b,*])
67 (=[b,i0])
68 )
69 )
70 (>[b]
71 (in[b,N])
72 (>[a,c,d]
73 (&
74 (in2[b,c,s])
75 (in3[a,b,d,*])
76 )
77 (&
78 (>[e]
79 (in3[d,a,e,+])
80 (in3[a,c,e,*])
81 )
82 (>[e]
83 (in3[a,c,e,*])
84 (in3[d,a,e,+])
85 )
86 )
87 )
88 )
89 )
90 )
91 )
92)
Listing 2: The definition of Natural Numbers in MPL.
3.3 Creation of Conjectures
Once definitions are formalized in MPL, GL begins the second stage of its pipeline: the automatic creation of conjectures. This process is fundamentally combinatorial, but it is constrained by several layers of control to manage the potential for explosive growth in the number of conjectures. The key challenge is to generate a rich set of potential theorems without being overwhelmed by syntactically valid but semantically useless statements.
A core strategy for managing this complexity is the use of a regularized theorem structure. Many mathematical theorems can be expressed, or easily transformed into, a linear chain of implications, such as $A\implies(B\implies(C\implies D))$. GL focuses on generating conjectures that fit this regularized form. This allows for a "weaving" process, where the system starts with a base LE and iteratively connects it with other building blocks using the implication operator.
To further suppress combinatorial explosion, GL applies several control mechanisms during this weaving process:
  1. 1.
    Normalization: After each new building block is added, all port names in the resulting LE are renamed according to a deterministic, canonical scheme. The system then considers all possible permutations of the building blocks within the nested implication structure and selects the one that is lexicographically smallest. This ensures that equivalent statements, such as $A\implies(B\implies C)$ and $B\implies(A\implies C)$, are represented by a single, unique conjecture.
  2. 2.
    Type-Based Connection Filtering: The process of connecting ports between two LEs is strongly restricted by the type system. In many domains, such as arithmetic, the system can be configured to only permit connections between ports that have identical definition set structures.
  3. 3.
    Topological Filtering: Users can define rules to exclude conjectures with certain structural properties. A common filter, for example, is one that discards any conjecture containing a circular dependency, where the output of one LE is indirectly used as an input to itself.
Through this combination of regularization, normalization, and filtering, GL can reduce the number of conjectures for a rich theory like Peano arithmetic from a potential billions to a manageable set of a few hundred.
3.4 CE Filter
Goal.
Before proof search, GL applies a deterministic CE filter to discard conjectures that are incompatible with small arithmetic tables.
Organization.
The CE filter reuses the same logic-block (LB) execution principle as the rest of GL, with a flat distribution: each conjecture is handled independently by its own LB. (A full description of LBs is deferred to the section dedicated to the prover.)
Inputs.
Each LB receives (i) a single conjecture and (ii) a finite set of simple ground facts, such as small tables for addition and multiplication (and related arithmetic facts). These tables serve as building blocks for CEs and the reference against which contradictions are tested.
Operation.
Within its LB, the filter incrementally weaves facts toward the conjecture’s premise using a simple peek-and-prune strategy: at each step it looks ahead just enough to decide whether to continue or to cut a branch. When a newly obtained fact contradicts one of the provided table facts, the conjecture is rejected immediately and the LB responsible for it is halted to conserve resources. Conjectures that survive this screening are forwarded to the prover.
Role in the pipeline.
The CE filter is a pre-proof triage mechanism: it removes many spurious conjectures early while leaving surviving conjectures unchanged for the prover. Its cost is small relative to conjecture generation and does not dominate memory or run time (RT), yet it substantially reduces downstream load.
3.5 Compilation
After a manageable set of conjectures has been generated, the system proceeds to the compilation phase. This phase translates the abstract logical structure of each conjecture into a concrete configuration for the grid of LBs. The process involves mapping, disintegration, and resource sharing.
3.5.1 Mapping to the Logic Grid
Each conjecture is mapped onto a chain of LBs. For an implication, the building blocks of its premise are assigned to a sequence of nodes. For example, the premise of the theorem $\forall n\in\mathbb{N}:0+n=n+0$ consists of two LEs: the definition of natural numbers and the statement $0+n$. These are mapped to two separate LBs connected in a directed path, where the output of the first is forwarded to the second, as shown in Figure 2. The conclusion of the implication is not mapped to a block; instead, it becomes the target output that the final block in the chain must produce.
Figure 2: Mapping a conjecture premise onto a sequence of LBs.
3.5.2 Disintegration
Once a LE is assigned to a block, the compilation process disintegrates it into its fundamental parts. Each sub-expression, such as a named function or a built-in operator, is extracted and stored in the block’s local memory. An implication within a definition is not disintegrated further but is stored in the block’s hash table to be used as a rule during execution. This "primordial soup" of logical components allows the block to recombine them during the execution phase to generate new expressions. For instance, an existence statement like $\exists y:P(x,y)$ is broken down, and a new, locally named port is created for the bound variable, enabling the block to reason about specific instances.
3.5.3 Resource Sharing via Tree Structuring
To optimize resource usage when compiling thousands of conjectures simultaneously, the system organizes them into a shared, tree-like structure. If multiple conjectures share a common prefix in their implication chain (e.g., they all rely on the definition of natural numbers), that shared prefix is mapped to a single set of LBs. The output of this shared root is then fanned out to different child branches representing the unique parts of each conjecture, as depicted in Figure 3. This ensures that each common logical component is computed only once, drastically reducing redundant calculations across the entire system.
Figure 3: Shared tree structure for multiple conjectures.
3.6 Processor
The final stage of the pipeline is the execution of the compiled conjectures on the GL processor grid. This phase is where the actual proof search occurs, driven by the asynchronous, parallel operation of the individual LBs.
3.6.1 Overview of Execution
Once LBs are compiled and configured, the execution sequence begins. Each block operates independently, without knowledge of or communication with other blocks during its computation cycle. All inter-block communication happens between cycles via the mail system. Before an execution, a block reads its incoming mail, and after, it sends out its outgoing mail. If a block proves a theorem that is valid for all blocks, this new theorem is provided to the GL system itself and broadcast to the entire network.
3.6.2 Core Execution Procedure
The core execution procedure within a single block is a simple but powerful process that forms the basis of all reasoning activities in GL. A single LB operates on three kinds of data:
  • Internally produced logical expressions.
  • Externally produced logical expressions received from parent nodes.
  • A local hash table containing implication rules derived from definitions.
The block’s main task is to generate a burst of hash requests by combining its internal and external expressions. To manage the combinatorial complexity of this task, the execution process is implemented as a variant of the peek-and-prune family of algorithms. The block weaves together combinations of expressions step-by-step, and after each extension, it checks if the resulting sequence exists as a subsequence in a prefix in its hash table. If not, that entire branch of computation is abandoned.
A further crucial optimization is the constraint that every hash key must contain at least one of the block’s locally generated LEs. This guarantees that any given inference can only be computed in one specific block, allowing for true distribution of RT usage across the different nodes of a theorem’s dependency tree.
3.6.3 Reasoning Contexts
This core execution procedure is versatile and can be deployed in different contexts to achieve various reasoning goals. The context determines how the output of the procedure is interpreted. For a proof by contradiction, the system uses the following approach: Assume a conjecture $A\implies(B\implies(C\implies D))$. This is mapped to a processing chain $A\to B\to C$. To prove it by contradiction, the system attaches the negated conclusion, $\neg D$, as a fourth block, creating the chain $A\to B\to C\to\neg D$. This final block is "primed" to detect a contradiction. If, during its execution, the $\neg D$ block generates or receives an expression that is the logical opposite of an expression it already holds (e.g., it derives $P(x)$ when it already possesses $\neg P(x)$), it signals a contradiction LB C. Upon receiving this signal, the LB C the original theorem, $A\implies(B\implies(C\implies D))$, as a new, globally valid theorem and broadcasts it to all other LBs. The same core execution procedure can be similarly adapted for other strategies, such as proof by induction or concrete symbolic calculations.
3.6.4 Conceptual Hardware Mapping
While GL is an abstract architecture that can be implemented in software, its design is heavily influenced by hardware co-design principles, making it particularly well-suited for a physical realization as an Application-Specific Integrated Circuit (ASIC). This section describes a sample hardware implementation based on this co-design philosophy.
A top-level view of a GL ASIC would comprise $N$ identical LBs, each paired with its own local SRAM macro. These blocks are interconnected by a shared on-chip bus or Network-on-Chip (NoC) supporting standard protocols like ARM AMBA AXI4. This bus must handle two distinct types of traffic for the mail system: low-throughput 32-bit "flag" messages for control signals (e.g., via AXI4-Lite) and variable-length "string" messages for logical expressions (e.g., via AXI4-Stream).
The internal architecture of each LB is a direct hardware mapping of the data structures required for symbolic reasoning. The hash-table-based string memory, for example, is realized using three distinct SRAM regions:
  • Bucket SRAM: A simple array of 32-bit pointers, where each entry points to the head of a collision chain in the HashNode region.
  • HashNode SRAM: A more complex, wider SRAM storing the nodes of the collision chains. Each entry holds pointers to the key string, the value string, and the next node in the chain.
  • String Data Slab: A large, byte-addressable SRAM region that stores the actual null-terminated strings for all keys and values used in the system.
Similarly, the asynchronous mail system’s inbox and outbox are implemented as hardware FIFOs, which decouple the internal computation of the block from the bus communication.
Finally, the system requires a compile-time configuration phase. Before execution, a centralized configuration master (such as an embedded CPU) uses the on-chip bus to pre-fill each LB’s local SRAMs with the necessary data to bootstrap the reasoning process. This includes the compiled routing tables (parent/child connections for mail), the Flag Table that maps flag IDs to RT actions, and any initial logical expressions or axioms required by the specific mathematical domain being explored. This pre-fill phase ensures that the static topology and initial state of the logic grid are correctly established before the first execution cycle begins.
3.7 Verifiable HTML Output
A core principle of GL is that its reasoning process must be transparent and its results independently verifiable. To this end, upon the completion of an execution run, GL generates a set of interlinked HTML documents that form a complete, human-readable proof graph. This output serves not only as the final record of the system’s discoveries but also as a powerful tool for debugging and visualization.
3.7.1 The Proof Index
The entry point to the generated proof graph is the index.html file. This page functions as a dynamic table of contents, listing all of the theorems that were successfully proven during the execution run. As shown in Figure 4, each entry in the index provides two representations of the theorem:
  1. 1.
    The formal, unabbreviated MPL expression that serves as the theorem’s unique identifier within the system.
  2. 2.
    A human-readable, algebraic simplification of the theorem, which provides an intuitive understanding of the proven statement.
Each theorem in the index is hyperlinked to its corresponding "chapter" page, which contains the detailed, step-by-step proof.
Figure 4: A screenshot of the main index.html page, showing the list of proven theorems with both their formal MPL and simplified algebraic representations.
3.7.2 The Interactive Proof Graph
Each "chapter" is a self-contained HTML page that details the proof of a single theorem. The page presents the sequence of LEs that were derived to reach the final conclusion. Each LE has its corresponding justification right next to it. The key feature of this output is its interactivity, which makes the proof fully auditable.
As illustrated in Figure 5, every step in the proof is annotated with a justification, such as "implication," "disintegration," or "recursion." Crucially, each justification is hyperlinked to the exact premise or previously proven theorem that was used to derive the current expression. A user can click on any step in the proof and be taken directly to the antecedent rule in the proof graph. This creates a fully traversable dependency chain, allowing a user to trace the reasoning for any theorem all the way back to the initial, foundational axioms. The goal of the proof is highlighted, and right-clicking on any MPL expression reveals a formatted, indented view for easier analysis.
This hyperlinked structure transforms a static list of statements into a dynamic and interactive proof graph. It allows for a level of scrutiny that is essential for verifying the system’s correctness. A researcher does not need to trust the internal workings of the GL system; the validity of its output can be confirmed simply by navigating the generated HTML files and verifying each logical step against its explicit justification. This commitment to transparent and verifiable output is fundamental to the design of GL.
Figure 5: A screenshot of a proof chapter (e.g., chapter17.html), showing several inference steps.
Table 1: End-to-end RT by phase on commodity hardware (32 logical cores).
Phase Time (s) Share (%)
Conjecturer 250 87.1
CE filter 30 10.5
Prover 7 2.4
Total 287 100
4 Case Study: Peano Arithmetic
We evaluate GL on first-order Peano arithmetic on commodity hardware (Dell G16 7630, 32 logical cores, 32 GB RAM). The end-to-end wall clock is about five minutes. The three phases contribute as follows: the conjecturer takes roughly 250 seconds, the CE filter about 30 seconds, and the prover about 7 seconds. The prover now runs at interactive speed; overall RT is dominated by conjecture generation and CE screening.
Rationale and headroom.
We deliberately defer aggressive micro-optimizations of the Peano pipeline. Many techniques can overfit to Peano’s shape and risk being useless or harmful on larger theories. Instead, we will pursue optimizations that transfer to Elementary Number Theory (ENT) before investing in phase-specific tuning. This preserves external validity while leaving substantial headroom to reduce the end-to-end RT.
What the numbers say.
  • Bottleneck is the conjecturer (about 87%). Transfer from Python to C++ is the easiest way to drive the current RT down due to stronger multithreading in the latter language. In practice, accelerating the conjecturer has never been a problem so far, taking the smallest amount of development time compared with the prover.
  • CE filter is modest but meaningful (about 10%). ENT-oriented speedups include batched evaluation, RT-friendly reformulation of conjectures, and reducing the amount of simple facts provided to the filter.
  • Prover is no longer the limit (about 2%). We expect a significant blow-up of prover RT for ENT. Ultimately, it depends on whether most wrong conjectures can be filtered by CEs. If, after all layers of aggressive filtering, there remain, for example, 400 mostly correct conjectures, RT should stay under control even with trichotomy.
Takeaway. With a prover RT of approximately 7 seconds and peak memory below a gigabyte, GL’s current bottleneck is conjecture generation. We will drive the remaining RT down during the transition to ENT, prioritizing optimizations (prover runs with tiered parametrization, batched CE tiers) that generalize to larger theories.
The present RT distribution should be read as a snapshot of the current implementation rather than a stable characteristic of GL; as the system and target theories evolve, the relative weights may shift in ways that are hard to predict, though the prover will likely remain the principal bottleneck. Given the many yet-unexplored but plausible optimization avenues, we are moderately optimistic and eager to engage.
5 Related Work and Contributions
The field of automated reasoning has seen significant advancements from two primary directions: probabilistic models and interactive proof assistants. GL positions itself as a third paradigm, offering a deterministic and massively parallel approach to theorem discovery. This section situates GL in relation to prior work and clarifies its unique contributions.
LLMs
LLMs have demonstrated impressive capabilities in pattern matching and the reformulation of existing knowledge. They can often solve routine mathematical exercises and rephrase known proofs in a human-like manner. However, their underlying architecture is probabilistic, which presents a fundamental limitation for rigorous, multi-step deductive reasoning. LLMs do not possess an intrinsic model of logical truth; consequently, they struggle to generate novel, non-trivial proofs and are susceptible to producing plausible-sounding but logically flawed arguments ("hallucinations"). In contrast, GL is a deterministic system where every inference is a consequence of axiomatic definitions, making its outputs inherently verifiable. Recent neural approaches have started to make headway in formal theorem proving – for example, a 2025 open-source transformer model achieved about 60% success on a standard benchmark by training on millions of generated proofs [8] – but such data-intensive methods are orthogonal and complementary to GL’s logic-first architecture.
Interactive and Semi-Automated Proof Assistants.
Systems such as Lean, Rocq/Coq, and Isabelle/HOL provide expressive foundations for formalizing mathematics and verifying proofs in a small, trusted kernel [6]. Although their traditional workflow is interactive—users decompose goals, invoke tactics, and supply key intermediate lemmas—modern ecosystems expose substantial automation. Isabelle’s Sledgehammer tool, for example, exports the current goal together with a heuristic relevance slice of the local theory context to a portfolio of external automated theorem provers (ATPs) and SMT solvers (e.g., E, Vampire, Z3, cvc5); when one succeeds, Isabelle attempts proof reconstruction automatically [3]. Lean and Coq likewise support tactic automation and scripted, batch replay of large libraries. In short, contemporary ITPs can discharge many individual goals automatically once the goals and supporting libraries are in place. Coq’s ecosystem has recently incorporated “hammer” plugins (e.g. CoqHammer) that automatically attempt to prove goals by translating them to first-order logic and calling external provers [4] , often discharging nontrivial goals without manual guidance. Automated first-order provers like E have achieved remarkable success in solving individual conjectures, consistently ranking among the top systems in the annual CASC competitions[10]. Another premier first-order prover is Vampire, which has dominated many competition divisions – since 1999 it has won over 50 category titles in the “world cup” of theorem provers [7] – thanks to its highly optimized saturation engine. SMT solvers (e.g. Z3) excel at deciding logical formulas in rich theories automatically [5] , and they have become integral to many verification workflows (often being invoked behind the scenes by proof assistant tactics to discharge specific subgoals). The newest SMT solvers, such as cvc5, extend the capabilities of their predecessors (e.g. CVC4 and Z3) with support for a diverse range of theories and advanced features like higher-order reasoning and syntax-guided synthesis [2].
GL vs. ITP automation. GL differs not by rejecting such automation but by shifting the unit of work: instead of waiting for a human to pose a single conjecture, GL compiles a base set of definitions and then systematically explores their deductive closure, generating large families of conjectures and attempting proofs for all of them in parallel. Every inference step is a deterministic hash-prefix match recorded in an auditable proof graph, enabling bulk, post hoc verification outside the proving engine.
5.1 Contributions
The GL architecture introduces several novel features to the field of automated reasoning.
Definition-Initiated Bulk Theorem Generation. Where mainstream proof assistants are typically used in a goal-directed manner (a user states a conjecture and applies interactive and automated tactics—possibly calling out to external ATP/SMT tools—to discharge it), GL begins from a curated definition set. It then enumerates normalized implication templates, weaves well-typed combinations, and launches a distributed proof search over all resulting conjectures simultaneously. This enables automatic discovery of families of theorems whose statements were not explicitly posed by a user.
Deterministic and Verifiable by Construction
Every reasoning step in GL is a symbolic manipulation based on a hash-table lookup. The final output is a complete, auditable proof graph where each inference is explicitly linked to a prior definition or a previously proven theorem. This provides a level of trust and verifiability that is not available in probabilistic systems. The correctness of a proof can be independently checked by a simple, external script, without needing to trust the internal machinery of GL itself.
Massively Parallel Architecture
The core of GL is its distributed architecture, where the reasoning workload is divided among thousands or millions of independent LBs. Each block operates asynchronously, communicating only between cycles. This design allows GL to tackle computationally intensive search problems by leveraging modern parallel computing infrastructure, such as multi-core ASICs or cloud-based server clusters.
Hardware-Software Co-Design
GL is conceived as a complete system, with the algorithms and architecture designed to complement each other. This hardware-aware approach ensures that the abstract reasoning process can be efficiently mapped onto a physical or virtualized processor, making it a scalable and practical solution for large-scale mathematical exploration.
6 Future Work, Roadmap and Potential Impact
The successful application of GL to Peano arithmetic serves as a crucial proof of concept, validating the core architecture and its approach to automated reasoning. However, this is merely the first step. The ultimate vision for GL is to tackle progressively more complex mathematical domains. This section outlines the next major milestone for the project and the concrete steps required to achieve it.
6.1 Next Milestone: ENT
The next grand challenge for GL is the automatic formulation of ENT. The goal is to provide the system with the necessary foundational definitions and have it autonomously generate a complete, verifiable proof graph for all major theorems up to and including the Fundamental Theorem of Arithmetic (FTA). Achieving this would represent a significant leap from the current baseline, demonstrating that the GL architecture is not limited to a single, simple theory but is a general-purpose engine for mathematical exploration.
6.2 The Path to Scalability and Performance
The roadmap to achieving the required performance involves a three-stage optimization and scaling plan:
  1. 1.
    Codebase Migration: To further enhance performance, the entire prototype will be migrated from Python to C++. This will provide significant speed improvements due to C++’s lower-level memory management and more efficient execution.
  2. 2.
    Massive Parallelization: The final and most critical step is to fully realize the parallel nature of the GL architecture. It will be adapted to run on massively parallel cloud infrastructure, distributing the workload across thousands of spot instances on platforms like Amazon Web Services (AWS) or Google Cloud Platform (GCP).
6.3 Architectural Enhancements for Advanced Reasoning
Beyond performance, tackling ENT requires several key enhancements to the core capabilities of the GL engine itself. The following architectural upgrades are planned:
Definition Integration
Currently, GL is proficient at disintegrating existing LEs into their constituent parts for use in proofs. The system must be extended with the inverse capability: the ability to integrate or assemble new, complex LEs from available building blocks, such as creating a LE for inequality from the LE of sum and concept of existence.
Introduction of Negation
The theorem weaving process will be enhanced to include negation. This will allow the system to automatically formulate a much richer set of conjectures, including those of the form $\neg A\implies B$ or $A\implies\neg B$.
Implementation of Trichotomy
The proof process will be augmented with the law of trichotomy, allowing the system to reason by cases (e.g., for any two numbers $a$ and $b$, either $a<b$, $a=b$, or $a>b$). This is an essential tool for many proofs in number theory.
By systematically executing this roadmap of performance optimizations and architectural enhancements, we anticipate that GL will be equipped to cover the domain of ENT within a reasonable RT, marking a major step towards the goal of fully automated, large-scale mathematical discovery.
6.4 LLM Deployment
LLM Interface & Feedback Loop (post‑ENT). After the ENT milestone, we will expose tool calls so an LLM can (i) submit MPL definitions/conjectures, (ii) invoke GL runs, and (iii) inspect derived theorems and proof graphs. In the reverse direction, GL becomes an LLM reasoning booster: when the model confronts a mathematical (and eventually any MPL‑encodable) reasoning task, it can translate that task to MPL and ask GL to deterministically study its deductive consequences, feeding validated results back into the dialogue. GL remains an external proof engine—not a Python library import—so the LLM/GL handshake mirrors, but does not collapse into, standard Python execution flows.
7 Potential Impact (Speculative)
The architectural path sketched in this work suggests several long‑horizon applications if the performance milestones are met.
  1. 1.
    Large‑scale axiom exploration. Million‑fold throughput would permit exhaustive closure over richer theories (ENT, Algebra, selected fragments of Analysis), enabling automated conjecture mining across domains.
  2. 2.
    Deterministic back‑end for AI agents. LLMs could emit conjecture invariants or lemmas and defer all symbolic checking to GL, yielding verifiable reasoning traces.
  3. 3.
    Formal assurance at design scale. Embedding GL cores in EDA/HW verification loops could shrink overnight proof farms to on‑device checks, improving safety‑critical certification.
  4. 4.
    Cross‑theory discovery. Uniform normalisation across imported definition sets might surface unexpected links (e.g., arithmetic/algebra correspondences) that merit human follow‑up.
7.1 LLMs Integration
Bidirectional LLM–GL Synergy. Today’s LLM coding assistants improve developer productivity by writing Python that a trusted interpreter executes. We aim for an analogous—but higher‑stakes—loop in formal reasoning: LLMs learn the MPL used by GL, submit definitional fragments or conjectures, and invoke GL’s deterministic, definition‑first exploration engine to mine consequences and produce fully traceable proof graphs. In return, GL functions as an LLM enhancer: whenever the model faces a mathematical (and potentially broader structured) reasoning task, it can externalize that task into MPL, let GL study it rigorously, and incorporate the resulting validated theorems—or contradictions—back into its natural‑language deliberation. Because GL records provenance for every derived step and was designed for scalable, auditable theorem generation from definitions, this closed loop could add a layer of trust and rigor to LLM‑mediated mathematical workflows well beyond what probabilistic text generation alone can provide. (Exploratory; not yet implemented.)
8 Conclusion
Generative Logic (GL) set out to test a simple question: Can a deterministic, definition‑centric architecture automatically discover and verify non‑trivial theorems from first principles? In this paper we demonstrated an end‑to‑end baseline: starting from a formalised Peano axiom set in MPL, GL generated conjectures and constructed verifiable proofs for the six foundational arithmetic laws, emitting a fully navigable HTML proof graph for audit.
Three architectural properties were validated empirically. First, the hash‑table inference model enables transparent, stepwise replay of every derivation. Second, compilation of definitions into distributed logic blocks cleanly separates symbolic content from execution, supporting parallel scaling. Third, the definition‑first workflow shifts human effort to axiom design; once definitions compile, proofs follow automatically (Section 2).
The current implementation remains a point‑zero prototype: single‑core Python, string‑based data structures, and modest performance headroom. Section 6 outlines the path to scale—C++ realisation, richer theorem weaving (negation, trichotomy)—on the way to ENT and beyond.
We view GL not as a replacement for human mathematicians but as a new instrument: a deterministic engine that can explore the deductive consequences of axioms at machine speed, extending the reach of formal reasoning.
These directions are intentionally forward‑looking; we present the Peano baseline in this paper to enable the community to evaluate, replicate, and extend the approach.
9 List of Abbreviations
AI
Artificial Intelligence
ASIC
Application-Specific Integrated Circuit
GL
Generative Logic
LB
Logic Block
RT
Run Time
HW
Hardware
ITP
Interactive Theorem Prover
LE
Logical Entity
LLM
Large Language Model
MPL
Mathematical Programming Language
SW
Software
ENT
Elementary Number Theory
FTA
Fundamental Theorem of Arithmetic
AWS
Amazon Web Service
GCP
Google Cloud Platform
CE
Counterexample