### v3

# Generative Logic: A New Computer Architecture for Deterministic Reasoning and Knowledge Generation

Abstract
We present Generative Logic (GL), a deterministic architecture that starts from user-supplied axiomatic definitions, written in a minimalist Mathematical Programming Language (MPL), and systematically explores a configurable region of their deductive neighborhood. A defining feature of the architecture is its unified hash-based inference engine, which executes both algebraic manipulations and deterministic logical transformations. Definitions are compiled into a distributed grid of simple Logic Blocks (LBs) that exchange messages; whenever the premises of an inference rule unify, a new fact is emitted with full provenance to its sources, yielding replayable, auditable proof graphs. Experimental validation is performed on Elementary Number Theory (ENT) utilizing a batched execution strategy. Starting from foundational axioms and definitions, the system first develops first-order Peano arithmetic, which is subsequently applied to autonomously derive and prove Gauss’s summation formula as a main result. To manage combinatorial explosion, GL algorithmically enumerates conjectures and applies normalization, type constraints, and counterexample (CE) filtering. On commodity hardware, an end-to-end run completes in under 7 minutes. Generated proofs export as navigable HTML so that every inference step can be inspected independently. We outline a hardware-software co-design path toward massively parallel realizations and describe future integration with large language models (LLMs) for auto-formalization and conjecture seeding. The Python, C++, and MPL code to reproduce these experiments, along with the full proof graphs in HTML as well as machine-readable text format, are available in the project’s GitHub repository at github.com/Generative-Logic/GL commit 1771330 and are permanently archived at doi:10.5281/zenodo.17206386.
1 Introduction
The pursuit of automated mathematical reasoning currently stands at a crossroads, defined by two distinct paradigms. On one side, Large Language Models (LLMs) have demonstrated a remarkable capacity for solving routine problems and reformulating existing knowledge, yet their probabilistic nature limits their reliability in discovering novel, non-trivial proofs [11]. On the other side, interactive proof assistants such as Lean and Coq offer formal guarantees but remain fundamentally manual endeavors, requiring extensive human expertise and guidance [9]. This leaves the vast landscape of deep, creative mathematical exploration largely inaccessible to full automation.
This paper introduces Generative Logic (GL), a novel computer architecture conceived to bridge this gap. GL is designed to automate deep, definition-based mathematical reasoning by shifting the paradigm from human-assisted, single-theorem proving to the fully automatic generation and verification of entire families of theorems. At its core, the system takes a set of formal axioms and algorithmically explores the constrained deductive space they define. This approach enables GL to systematically generate conjectures, produce machine-verifiable proofs, and iterate through complex logical chains at computational speeds.
The implications of such a system are significant. By accelerating the cycle of discovery and verification, GL has the potential to act as a catalyst for mathematical research and to shorten the path from abstract breakthroughs to downstream innovations in physics, engineering, and computer science. Furthermore, we envision GL serving as a deterministic reasoning core that can be integrated into probabilistic Artificial Intelligence (AI) workflows, providing a source of verifiable truth to enhance the capabilities of next-generation AI systems – a vision shared by recent work calling for the fusion of LLMs with formal methods to build more trustworthy AI [12]. In essence, just as the electronic calculator democratized arithmetic, GL aims to democratize high-level mathematical reasoning.
This paper presents the foundational principles, architecture, and early results of GL. We detail the system’s massively parallel, hardware-aware proof engine, its custom Mathematical Programming Language (MPL) for formalizing definitions, and its unique execution model. A case study in Elementary Number Theory (ENT) demonstrates the system’s ability to autonomously derive and prove foundational theorems, validating the architecture and its approach to automated discovery.
In this light, GL can be understood through an analogy with biological development. A living organism does not store an explicit blueprint of every cell, tissue, and organ it will produce. Instead, a compact set of genetic instructions — DNA — encodes a small collection of foundational rules, and the machinery of the cell does the rest: molecules interact, differentiate, and self-organize into structures of extraordinary complexity. Critically, this unfolding is not unbounded. Gene regulation, energy budgets, and environmental signals determine which developmental paths are expressed in a given context; the genome encodes far more potential than any single organism realizes. GL operates on the same principle, but in the domain of logic. A minimal set of axiomatic definitions, written in MPL, plays the role of the genome: a compact, high-density encoding of foundational truth. The GL architecture then acts as the developmental machinery — a deterministic engine that explores the deductive consequences those definitions imply, with per-batch configuration files playing a role analogous to gene regulation: they do not alter the logical content of the definitions, but they determine which region of the deductive space is explored in a given run. Just as no biologist needs to manually specify every protein interaction for an organism to grow, no mathematician needs to manually state every theorem for GL to discover it. The axioms are planted; the proof graph grows. As computational resources scale, constraints can be progressively relaxed, allowing broader exploration from the same definitional base. This shift — from explicit enumeration to generative unfolding — is the central architectural commitment of GL, and it is what distinguishes it from both interactive proof assistants, which require human-directed goal-setting, and from probabilistic language models, which approximate reasoning without grounding it in axiomatic truth.
2 The GL Workflow
The GL system transforms a set of formal definitions into a web of proven theorems through a deterministic, multi-stage pipeline. This process is designed to be exhaustive, shifting the primary human contribution to the initial formulation of clear and consistent definitions. The workflow, outlined below, mirrors a software development cycle where errors in definitions can lead to "compilation" failures or run-time contradictions, which are themselves valuable for debugging the axiomatic foundation. Optionally, simple facts expressed in MPL (e.g., small addition/multiplication tables) can be supplied to filter out wrong conjectures and to reduce resource consumption.
The GL pipeline is executed in discrete, ordered batches, where each batch corresponds to a self-contained mathematical domain or theory layer. This batched design serves a primary practical purpose: by demixing domains, it prevents the conjecturer and prover from operating over a conflated search space, which would combinatorially explode as the number of definitions grows. Instead, each batch operates on a focused, well-scoped slice of the theory, with its results becoming available as proven ground for subsequent batches. Each batch is governed by a dedicated configuration file, which exercises two key controls: it defines the search space of the conjecturer — specifying which definitions, operators, and structural templates are eligible for conjecture weaving — and it sets the operational depth of the prover, determining how far the inference engine will chase logical consequences before halting. This separation of concerns between batches, combined with per-batch configurability, is what allows GL to scale incrementally toward richer theories without sacrificing the determinism and auditability that define the architecture.
  1. 1.
    Definition Input. The process begins with a user, or an LLM assistant, providing a set of foundational axioms and definitions. These are formalized using MPL, a domain-specific language designed to express second-order predicate logic in a machine-readable format. The integrity of the entire reasoning process depends on the quality of this initial input.
  2. 2.
    Conjecture Generation. Using the provided definitions, GL automatically and combinatorially generates a large set of conjectures. To manage the combinatorial explosion inherent in this step, GL employs a "weaving" process on a regularized theorem structure—typically of the form $A\implies(B\implies(C\implies D))$—and applies aggressive normalization and filtering techniques to discard redundant or malformed conjectures. This phase is purely syntactic and does not attempt to evaluate the semantic value of the conjectures.
  3. 3.
    Counterexample (CE) Filtering. A deterministic pre-proof triage stage that, for each conjecture, reuses GL’s prover architecture to test consistency against small arithmetic tables. Facts are woven with a peek-and-prune strategy to generate new facts based on the conjecture under test; any contradiction triggers immediate rejection.
  4. 4.
    Compilation and Disintegration. The generated conjectures are then compiled and distributed across the grid of Logic Blocks (LBs). Each Logical Entity (LE) within a conjecture is assigned to a specific block, forming a distributed processing chain. Within each block, the relevant definitions are disintegrated into their constituent logical parts, which serve as the initial fuel for the proof search.
  5. 5.
    Iterative Proof Execution. With the system primed, GL begins a "flood" of asynchronous iterations to find proofs for all conjectures simultaneously. This core execution process is fundamentally symbolic and hash-table based. For a given implication, the premise acts as a hash request (the key) and the conclusion as the expected response (the value). During each computation cycle, a LB performs a burst of hash requests, combining its local expressions with incoming mail to probe the vast space of logical consequences.
    This process is deeply distributed; each LB operates independently on its local memory, and communication with other blocks only occurs between cycles. This design allows for massive parallelization across thousands or even millions of cores, turning a computationally intractable search into a manageable, parallel process.
  6. 6.
    Verifiable Output. Once the execution run is complete, GL generates a set of HTML files that serve as the final, verifiable output for visualization and debugging. These files contain a human-readable list of all proven theorems. Crucially, each step in every proof is hyperlinked to its justification—either an initial definition or a previously proven theorem—creating a fully auditable and transparent proof graph that can be independently verified.
The natural habitat of GL is a cloud of specialized ASIC chips — a massively parallel substrate where multitudes of Logic Blocks operate simultaneously and the architecture’s full potential is realized without constraint. For the time being, however, GL must survive in the bonsai / resource-constrained conditions of a single machine or a cloud fragment. Configuration files are precisely the instrument that makes this possible. By tightening the conjecturer’s search space and capping the prover’s operational depth, they allow available computational power to be traded against domain-specific knowledge: the less compute one has, the more the configuration must compensate by encoding structural guidance, type constraints, and targeted conjecture templates. Conversely, as hardware scales up, constraints can be relaxed and the system allowed to explore more freely. This mechanism is what tames a beast built for unlimited compute and lets it run, usefully, on a local machine or cloud fragment. The smallest but still meaningful deployment of GL is one where the user supplies not only definitions but all necessary lemmas alongside them, leaving GL the focused task of assembling these ingredients into a verified, navigable proof graph — sacrificing breadth of discovery in exchange for guaranteed tractability.
In the current implementation, simple facts such as addition and multiplication tables are supplied manually by the user. A near-future release will automate this step through a dedicated tool called the Incubator. The key insight is that each such fact — for example, that 3 × 4 = 12 — is itself a simple theorem, and as such can in principle be conjectured, filtered, and proven by GL’s own machinery. The Incubator will be guided by a dedicated configuration file that directs GL to conjecture and either prove or disprove ground-level arithmetic facts, effectively allowing the system to bootstrap its own fact base. However, GL currently lacks case differentiation — the ability to reason by cases such as trichotomy — which is a prerequisite for constructing proofs of this kind. Once case differentiation is enabled, the Incubator will follow as a natural consequence, closing the loop and making the pipeline fully self-sufficient from definitions down to ground facts.
3 System Architecture
The GL workflow is supported by a unique hardware-aware architecture designed for massive parallelism and deterministic symbolic manipulation. The system is not a monolithic processor but a distributed network of simple, independent processing elements that collectively perform complex reasoning tasks. This design can be realized physically as an Application-Specific Integrated Circuit (ASIC) or implemented in software on cloud infrastructure or multi-core servers.
3.1 Architecture Overview
The core of GL is a symbolic reasoning engine. Unlike traditional theorem provers that rely on heuristic search algorithms, GL treats logical inference as a memory-access problem. In this paradigm, an implication is modeled as a key-value pair in a distributed hash table. For an implication such as $(A\land B)\implies C$:
  • The premise, $(A\land B)$, serves as the hash key.
  • The conclusion, $C$, serves as the hash value.
Reasoning is thus transformed into a process of formulating hash requests from known logical expressions and, upon a successful lookup, receiving a new, valid expression as the response. Both the keys and values are represented as plain text strings, underscoring the symbolic nature of the architecture.
This engine is realized as a deeply distributed system of independent nodes, or LBs, which are connected through configurable, directed paths. Each LB possesses its own local memory and executes its core operation—a burst of hash requests—asynchronously. Newly generated expressions are passed to other nodes according to a pre-compiled connection scheme. This block-based architecture, where each node operates independently, is the key to the system’s scalability. It allows the computational load to be parallelized across thousands or even millions of cores, turning a search problem that would be intractable on a single processor into a manageable, massively parallel task.
Consider the following example. Full Implication:
$(n>m\land m>2)\implies(n>2)$
Premise:
$n>m\land m>2$
Conclusion:
$n>2$
Figure 1: An example for connection of two LBs.
Figure 1 visualizes in a simple way how two LBs can be connected.
3.2 MPL
The GL system requires a formal, machine-readable way to express mathematical definitions and axioms. To this end, we developed MPL, a language based on a constrained subset of second-order predicate logic. MPL is designed to be expressive enough to capture complex mathematical structures while remaining sufficiently simple for algorithmic analysis and compilation. This corresponds to the well-established logical framework of higher-order logic, as detailed in [1].
3.2.1 LEs and Ports
The fundamental unit in MPL is the LE, which represents a predicate that evaluates to a Boolean value. An LE can be conceptualized as a template function in a language like C++. Each LE has a set of "ports" through which it receives arguments. For example, the LE representing the statement $x>2$ has three ports: one for $x$, one for the constant $2$, and one for the operator >>.
Each port is defined by a definition set, which functions as a type system. For the LE $x>2$, the ports for $x$ and $2$ might have the definition set $\mathbb{N}$ (the set of natural numbers), while the port for >> would have the definition set $\mathcal{P}(\mathbb{N}\times\mathbb{N})$ (the set of binary relations on $\mathbb{N}$). This typed approach allows for strong compile-time checks on the validity of logical constructions.
3.2.2 Core Operators
New LEs are constructed from existing ones using a minimal set of three core operators, which are sufficient to express all of standard logic.
Negation (!)
The ! operator takes a single LE and inverts its Boolean output. For example, !(in[x,A]) corresponds to $x\notin A$. The ports and their definition sets remain unchanged.
Conjunction (&)
The (& A B) operator combines two LEs, $A$ and $B$, into a single LE that evaluates to true if and only if both $A$ and $B$ are true. Ports from $A$ and $B$ can be merged by assigning them the same reference name. For instance, in (& (in2[n,m,>]) (in2[m,q,<])), the port ‘m‘ is shared, effectively unifying the corresponding definition sets and creating a dependency between the two original LEs.
Implication (>>)
The (>[vars](A)(B)) operator is the most complex. If the variable list ‘[vars]‘ is empty, it represents logical implication, equivalent to $\neg A\lor B$. If the list is non-empty, such as in (>[n](A)(B)), it represents the universal quantifier "for all". This operator binds the variable ‘n‘ within the scope of the expression, asserting that for all possible values of ‘n‘ (as defined by its port’s definition set), the implication $\neg A(n)\lor B(n)$ holds.
The existential quantifier, $\exists$, is not needed as a primitive operator, as it can be constructed from negation and the universal quantifier: $\exists x:P(x)\equiv\neg(\forall x:\neg P(x))$.
3.2.3 Type System and Recursion Constraint
The MPL compiler enforces a strict type system when merging ports. A critical rule is the prevention of infinite recursive definitions. A definition set for a port cannot be unified with a set that contains it. For example, a port of type $X$ cannot be merged with a port of type $\mathcal{P}(X\times Y)$, as this would imply that the set $X$ contains a structure defined by itself. This restriction, analogous to type safety in programming languages, prevents paradoxes and ensures that all definitions are well-founded, though it limits the direct expression of concepts such as ordinal numbers.
3.2.4 Unique Naming of Reusable LEs
Complex, reusable concepts can be defined once and given a name. The listing below shows how the mathematical definition of a function $f:X\to Y$ is expressed in MPL. This LE, named ‘fXY‘, encapsulates the properties that for every element $x$ in the domain $X$, there exists a unique element $y$ in the codomain $Y$ with $f(x)=y$ .
1(fXY[f,X,Y]) :=
2(&
3 (>[x,y]
4 (in2[x,y,f])
5 (&
6 (in[x,X])
7 (in[y,Y])
8 )
9 )
10 (&
11 (>[x]
12 (in[x,X])
13 ! (>[y]
14 (in[y,Y])
15 ! (in2[x,y,f])
16 )
17 )
18 (>[x]
19 (in[x,X])
20 (>[y1,y2]
21 (&
22 (in2[x,y1,f])
23 (in2[x,y2,f])
24 )
25 (=[y1,y2])
26 )
27 )
28 )
29)
Listing 1: The definition of a function $f:X\to Y$ in MPL.
Another example of a foundational definition provided to the system, the Peano axioms for natural numbers, is shown below.
1(NaturalNumbers[N,i0,i1,s,+,*]) :=
2(&
3 (&
4 (in[i0,N])
5 (&
6 (in2[i0,i1,s])
7 (&
8 (fXY[s,N,N])
9 (&
10 (>[n]
11 (in[n,N])
12 !(in2[n,i0,s])
13 )
14 (>[m]
15 (in[m,N])
16 (>[n1,n2]
17 (&
18 (in2[n1,m,s])
19 (in2[n2,m,s])
20 )
21 (=[n1,n2])
22 )
23 )
24 )
25 )
26 )
27 )
28 (&
29 (&
30 (fXYZ[+,N,N,N])
31 (&
32 (>[a]
33 (in[a,N])
34 (>[b]
35 (in3[a,i0,b,+])
36 (=[a,b])
37 )
38 )
39 (>[b]
40 (in[b,N])
41 (>[a,c,d]
42 (&
43 (in2[b,c,s])
44 (in3[a,b,d,+])
45 )
46 (&
47 (>[e]
48 (in3[a,c,e,+])
49 (in2[d,e,s])
50 )
51 (>[e]
52 (in2[d,e,s])
53 (in3[a,c,e,+])
54 )
55 )
56 )
57 )
58 )
59 )
60 (&
61 (fXYZ[*,N,N,N])
62 (&
63 (>[a]
64 (in[a,N])
65 (>[b]
66 (in3[a,i0,b,*])
67 (=[b,i0])
68 )
69 )
70 (>[b]
71 (in[b,N])
72 (>[a,c,d]
73 (&
74 (in2[b,c,s])
75 (in3[a,b,d,*])
76 )
77 (&
78 (>[e]
79 (in3[d,a,e,+])
80 (in3[a,c,e,*])
81 )
82 (>[e]
83 (in3[a,c,e,*])
84 (in3[d,a,e,+])
85 )
86 )
87 )
88 )
89 )
90 )
91 )
92)
Listing 2: The definition of Natural Numbers in MPL.
3.3 Creation of Conjectures
Once definitions are formalized in MPL, GL begins the second stage of its pipeline: the automatic creation of conjectures. This process is fundamentally combinatorial, but it is constrained by several layers of control to manage the potential for explosive growth in the number of conjectures. The key challenge is to generate a rich set of potential theorems without being overwhelmed by syntactically valid but semantically useless statements.
A core strategy for managing this complexity is the use of a regularized theorem structure. Many mathematical theorems can be expressed, or easily transformed into, a linear chain of implications, such as $A\implies(B\implies(C\implies D))$. GL focuses on generating conjectures that fit this regularized form. This allows for a "weaving" process, where the system starts with a base LE and iteratively connects it with other building blocks using the implication operator.
To further suppress combinatorial explosion, GL applies several control mechanisms during this weaving process:
  1. 1.
    Normalization: After each new building block is added, all port names in the resulting LE are renamed according to a deterministic, canonical scheme. The system then considers all possible permutations of the building blocks within the nested implication structure and selects the one that is lexicographically smallest. This ensures that equivalent statements, such as $A\implies(B\implies C)$ and $B\implies(A\implies C)$, are represented by a single, unique conjecture.
  2. 2.
    Type-Based Connection Filtering: The process of connecting ports between two LEs is strongly restricted by the type system. In many domains, such as arithmetic, the system can be configured to only permit connections between ports that have identical definition set structures.
  3. 3.
    Topological Filtering: Users can define rules to exclude conjectures with certain structural properties. A common filter, for example, is one that discards any conjecture containing a circular dependency, where the output of one LE is indirectly used as an input to itself.
Through this combination of regularization, normalization, and filtering, GL can reduce the number of conjectures for a rich theory like Peano arithmetic from a potential billions to a manageable set of a few hundred.
3.4 CE Filter
The CE filter is as a rejection mechanism relative to the supplied ground-fact tables: if a contradiction is found, the conjecture is discarded. However, it is incomplete as a decision procedure: surviving the CE stage does not imply truth or provability, only compatibility with the tested finite facts.
Goal.
Before proof search, GL applies a deterministic CE filter to discard conjectures that are incompatible with small arithmetic tables.
Organization.
The CE filter reuses the same logic-block (LB) execution principle as the rest of GL, with a flat distribution: each conjecture is handled independently by its own LB. (A full description of LBs is deferred to the section dedicated to the prover.)
Inputs.
Each LB receives (i) a single conjecture and (ii) a finite set of simple ground facts, such as small tables for addition and multiplication (and related arithmetic facts). These tables serve as building blocks for CEs and the reference against which contradictions are tested.
Operation.
Within its LB, the filter incrementally weaves facts toward the conjecture’s premise using a simple peek-and-prune strategy: at each step it looks ahead just enough to decide whether to continue or to cut a branch. When a newly obtained fact contradicts one of the provided table facts, the conjecture is rejected immediately and the LB responsible for it is halted to conserve resources. Conjectures that survive this screening are forwarded to the prover.
Role in the pipeline.
The CE filter is a pre-proof triage mechanism: it removes many spurious conjectures early while leaving surviving conjectures unchanged for the prover. Its cost is small relative to conjecture generation and does not dominate memory or run time (RT), yet it substantially reduces downstream load.
Upcoming extension - Incubator.
In the current implementation, simple facts used by the CE filter are supplied manually by the user. A near-future release will automate this step through a dedicated tool called the Incubator. The key insight is that each such fact is itself a simple theorem, and as such can in principle be conjectured, filtered, and proven by GL’s own machinery, regardless of the domain under exploration. The Incubator will be guided by a dedicated configuration file that directs GL to conjecture and either prove or disprove ground-level facts, effectively allowing the system to bootstrap its own fact base. This tool becomes increasingly necessary as the complexity of the target domain grows: what counts as a "simple" fact in elementary arithmetic is a straightforward table entry, but in richer domains, the corresponding ground-level facts can themselves be non-trivial statements requiring structured derivation. Without automation, the manual curation of such facts quickly becomes a bottleneck. However, GL currently lacks case differentiation — the ability to reason by cases — which is a prerequisite for constructing proofs of this kind. Once case differentiation is enabled, the Incubator will follow as a natural consequence, closing the loop and making the pipeline fully self-sufficient from definitions down to ground facts, across any domain expressible in MPL.
3.5 Compilation
After a manageable set of conjectures has been generated, the system proceeds to the compilation phase. This phase translates the abstract logical structure of each conjecture into a concrete configuration for the grid of LBs. The process involves mapping, disintegration, and resource sharing.
3.5.1 Mapping to the Logic Grid
Each conjecture is mapped onto a chain of LBs. For an implication, the building blocks of its premise are assigned to a sequence of nodes. For example, the premise of the theorem $\forall n\in\mathbb{N}:0+n=n+0$ consists of two LEs: the definition of natural numbers and the statement $0+n$. These are mapped to two separate LBs connected in a directed path, where the output of the first is forwarded to the second, as shown in Figure 2. The conclusion of the implication is not mapped to a block; instead, it becomes the target output that the final block in the chain must produce.
Figure 2: Mapping a conjecture premise onto a sequence of LBs.
3.5.2 Disintegration
Once a LE is assigned to a block, the compilation process disintegrates it into its fundamental parts. Each sub-expression, such as a named function or a built-in operator, is extracted and stored in the block’s local memory. An implication within a definition is not disintegrated further but is stored in the block’s hash table to be used as a rule during execution. This "primordial soup" of logical components allows the block to recombine them during the execution phase to generate new expressions. For instance, an existence statement like $\exists y:P(x,y)$ is broken down, and a new, locally named port is created for the bound variable, enabling the block to reason about specific instances.
3.5.3 Resource Sharing via Tree Structuring
To optimize resource usage when compiling thousands of conjectures simultaneously, the system organizes them into a shared, tree-like structure. If multiple conjectures share a common prefix in their implication chain (e.g., they all rely on the definition of natural numbers), that shared prefix is mapped to a single set of LBs. The output of this shared root is then fanned out to different child branches representing the unique parts of each conjecture, as depicted in Figure 3. This ensures that each common logical component is computed only once, drastically reducing redundant calculations across the entire system.
Figure 3: Shared tree structure for multiple conjectures.
3.6 Processor
The final stage of the pipeline is the execution of the compiled conjectures on the GL processor grid. This phase is where the actual proof search occurs, driven by the asynchronous, parallel operation of the individual LBs.
3.6.1 Overview of Execution
Once LBs are compiled and configured, the execution sequence begins. Each block operates independently, without knowledge of or communication with other blocks during its computation cycle. All inter-block communication happens between cycles via the mail system. Before an execution, a block reads its incoming mail, and after, it sends out its outgoing mail. If a block proves a theorem that is valid for all blocks, this new theorem is provided to the GL system itself and broadcast to the entire network.
3.6.2 Core Execution Procedure
The core execution procedure within a single block is a simple but powerful process that forms the basis of all reasoning activities in GL. A single LB operates on three kinds of data:
  • Internally produced logical expressions.
  • Externally produced logical expressions received from parent nodes.
  • A local hash table containing implication rules derived from definitions.
The block’s main task is to generate a burst of hash requests by combining its internal and external expressions. To manage the combinatorial complexity of this task, the execution process is implemented as a variant of the peek-and-prune family of algorithms. The block weaves together combinations of expressions step-by-step, and after each extension, it checks if the resulting sequence exists as a subsequence in a prefix in its hash table. If not, that entire branch of computation is abandoned.
A further crucial optimization is the constraint that every hash key must contain at least one of the block’s locally generated LEs. This guarantees that any given inference can only be computed in one specific block, allowing for true distribution of RT usage across the different nodes of a theorem’s dependency tree.
3.6.3 Deterministic Logical Transformer
The same hash-based inference engine that drives proof search through the application of previously proven theorems — that is, algebraic manipulation — is also responsible for a fundamentally different but equally central task: the deterministic transformation of existing logical entities into new ones, a capability we term the Deterministic Logical Transformer. (The term ’transformer’ is used here in its classical sense — a mechanism that converts one structure into another — and bears no relation to the neural Transformer architecture of Vaswani et al.)
The transformation process consists of two complementary operations, both introduced earlier in the pipeline: disintegration and integration. Disintegration decomposes a compound logical entity into its constituent parts; integration reassembles available parts into new, valid compound entities. Crucially, both operations are governed by the same underlying artifact: the GL binary, the principal output of the compilation phase. The binary encodes exact instructions for how any compound logical structure may be disassembled and how its components may be recombined. The compiled representation is therefore not directional — the same map can be traversed in both ways, making disintegration and integration two faces of a single deterministic process rather than separate mechanisms.
Figure 4 illustrates this structure for the fold definition. The root node is of type existence, which expands into two children: interval (type and) and existence2 (type existence). Each of these expands further into its constituent elements. Nodes of type implication are terminal — they are not expanded but stored as inference rules in the local hash table of the corresponding Logic Block. Predicate leaves represent ground logical statements that serve directly as facts or hash keys during execution. The same graph that guides disintegration — breaking fold down to its atomic parts — equally guides integration, assembling a valid fold instance from available components by traversing the map in reverse.
The roles of the three node categories differ fundamentally between the two directions of traversal. During disintegration, and and existence nodes act as predicate suppliers: their constituent expressions are extracted and injected into the hash engine as known facts, fueling the inference process. Implication nodes, by contrast, are loaded directly as key-value pairs into the hash table, where they serve as reusable inference rules available to the engine at any point during execution. During integration, the roles shift. Implication nodes now become sub-theorems that must be actively proved — the hash engine is tasked with establishing them before the compound entity can be assembled. Meanwhile, and and existence nodes undergo a structural inversion: they are themselves transformed into implications, reframing the question of constructing a compound entity as a conditional one. This allows the hash engine to reason about whether the necessary components are present and, if so, to conclude that the compound entity is validly constructible. Disintegration and integration thus engage the same underlying machinery in opposite modes — one dissolving structure into fuel, the other building structure from proof.
Figure 4: Compiled structure of the fold definition. Existence and and-nodes expand further; implication nodes are terminal stop points; predicate leaves are ground facts. The same map is traversed in both directions by the Deterministic Logical Transformer.
3.6.4 Reasoning Contexts
This core execution procedure is versatile and can be deployed in different contexts to achieve various reasoning goals. The context determines how the output of the procedure is interpreted. For a proof by contradiction, the system uses the following approach: Assume a conjecture $A\implies(B\implies(C\implies D))$. This is mapped to a processing chain $A\to B\to C$. To prove it by contradiction, the system attaches the negated conclusion, $\neg D$, as a fourth block, creating the chain $A\to B\to C\to\neg D$. This final block is "primed" to detect a contradiction. If, during its execution, the $\neg D$ block generates or receives an expression that is the logical opposite of an expression it already holds (e.g., it derives $P(x)$ when it already possesses $\neg P(x)$), it signals a contradiction LB C. Upon receiving this signal, the LB C emits the original theorem, $A\implies(B\implies(C\implies D))$, as a new, globally valid theorem and broadcasts it to all other LBs. The same core execution procedure can be similarly adapted for other strategies, such as proof by induction or concrete symbolic calculations.
3.6.5 Conceptual Hardware Mapping
While GL is an abstract architecture that can be implemented in software, its design is heavily influenced by hardware co-design principles, making it particularly well-suited for a physical realization as an Application-Specific Integrated Circuit (ASIC). This section describes a sample hardware implementation based on this co-design philosophy.
A top-level view of a GL ASIC would comprise $N$ identical LBs, each paired with its own local SRAM macro. These blocks are interconnected by a shared on-chip bus or Network-on-Chip (NoC) supporting standard protocols like ARM AMBA AXI4. This bus must handle two distinct types of traffic for the mail system: low-throughput 32-bit "flag" messages for control signals (e.g., via AXI4-Lite) and variable-length "string" messages for logical expressions (e.g., via AXI4-Stream).
The internal architecture of each LB is a direct hardware mapping of the data structures required for symbolic reasoning. The hash-table-based string memory, for example, is realized using three distinct SRAM regions:
  • Bucket SRAM: A simple array of 32-bit pointers, where each entry points to the head of a collision chain in the HashNode region.
  • HashNode SRAM: A more complex, wider SRAM storing the nodes of the collision chains. Each entry holds pointers to the key string, the value string, and the next node in the chain.
  • String Data Slab: A large, byte-addressable SRAM region that stores the actual null-terminated strings for all keys and values used in the system.
Similarly, the asynchronous mail system’s inbox and outbox are implemented as hardware FIFOs, which decouple the internal computation of the block from the bus communication.
Finally, the system requires a compile-time configuration phase. Before execution, a centralized configuration master (such as an embedded CPU) uses the on-chip bus to pre-fill each LB’s local SRAMs with the necessary data to bootstrap the reasoning process. This includes the compiled routing tables (parent/child connections for mail), the Flag Table that maps flag IDs to RT actions, and any initial logical expressions or axioms required by the specific mathematical domain being explored. This pre-fill phase ensures that the static topology and initial state of the logic grid are correctly established before the first execution cycle begins.
3.7 Verifiable HTML Output
A core principle of GL is that its reasoning process must be transparent and its results independently verifiable. To this end, upon the completion of an execution run, GL generates a set of interlinked HTML documents that form a complete, human-readable proof graph. This output serves not only as the final record of the system’s discoveries but also as a powerful tool for debugging and visualization.
3.7.1 The Proof Index
The entry point to the generated proof graph is the index.html file. This page functions as a dynamic table of contents, listing all of the theorems that were successfully proven during the execution run. As shown in Figure 5, each entry in the index provides two representations of the theorem:
  1. 1.
    The formal, unabbreviated MPL expression that serves as the theorem’s unique identifier within the system.
  2. 2.
    A human-readable, algebraic simplification of the theorem, which provides an intuitive understanding of the proven statement.
Each theorem in the index is hyperlinked to its corresponding "chapter" page, which contains the detailed, step-by-step proof.
Figure 5: A screenshot of the main index.html page, showing the list of proven theorems with both their formal MPL and simplified algebraic representations.
3.7.2 The Interactive Proof Graph
Each "chapter" is a self-contained HTML page that details the proof of a single theorem. The page presents the sequence of LEs that were derived to reach the final conclusion. Each LE has its corresponding justification right next to it. The key feature of this output is its interactivity, which makes the proof fully auditable.
As illustrated in Figure 6, every step in the proof is annotated with a justification, such as "implication," "disintegration," or "recursion." Crucially, each justification is hyperlinked to the exact premise or previously proven theorem that was used to derive the current expression. A user can click on any step in the proof and be taken directly to the antecedent rule in the proof graph. This creates a fully traversable dependency chain, allowing a user to trace the reasoning for any theorem all the way back to the initial, foundational axioms. The goal of the proof is highlighted, and right-clicking on any MPL expression reveals a formatted, indented view for easier analysis.
This hyperlinked structure transforms a static list of statements into a dynamic and interactive proof graph. It allows for a level of scrutiny that is essential for verifying the system’s correctness. A researcher does not need to trust the internal workings of the GL system; the validity of its output can be confirmed simply by navigating the generated HTML files and verifying each logical step against its explicit justification. This commitment to transparent and verifiable output is fundamental to the design of GL.
Figure 6: A screenshot of a proof chapter (chapter81.html), showing several inference steps.
4 Case Study: Elementary Number Theory
To assess the GL system, we conducted a two-stage evaluation encompassing both Peano arithmetic and Gauss’s summation formula. The experiments were performed on standard consumer hardware (a Dell G16 7630 equipped with 32 logical cores and 32 GB of RAM). The complete execution finished in roughly 407 seconds, or just under seven minutes. A detailed breakdown of the time spent in each phase across the two batches is provided in Table 1. Notably, rather than the standard fractional form, GL autonomously derives the division-free variant of Gauss’s formula, expressing the product of $n$ and its successor as twice the sum:
$n(n+1)=2\sum_{i=1}^{n}i$ (1)
Batch 1: Peano Arithmetic
The conjecturer produces 406 candidates in 21 seconds. The CE filter processes all 406 conjectures in 14.6 seconds, passing 55 survivors to the prover. The prover runs 17 iterations (2 warm-up, 15 main) and finishes in 24.6 seconds, establishing 32 theorems. These cover the foundational arithmetic laws of Peano arithmetic, including commutativity and associativity of addition and multiplication, and distributivity. The 64 theorem entries written to file represent mirrored pairs — each theorem together with its symmetric counterpart. The total runtime of the native executable for this batch is 39.4 seconds.
Batch 2: Gauss Summation
The Gauss batch begins by generating 363 conjectures in 106 seconds — the larger conjecture creation time reflecting the richer definition set, which now includes intervals, sequences, limit constructs, and the fold operator. Before the prover starts, the anchor connection mechanism automatically attaches the Gauss anchor to the Peano anchor, injecting the implication AnchorGauss $\Rightarrow$ AnchorPeano into the theorem list. This ensures that all 64 theorems proven in Batch 1 are available as ground knowledge for the Gauss prover.
The CE filter runs for 192.5 seconds — the dominant cost in this batch — passing 29 conjectures to the prover. At the start of Phase 2, the prover broadcasts the 64 inherited Peano theorems across all 70 active memory blocks, making prior knowledge immediately available to every Logic Block. The prover then runs 30 iterations (2 warm-up, 28 main) and finishes in 47.1 seconds, proving 10 theorems including the flagship result: Gauss’s summation formula. The total runtime of the native executable for this batch is 239.7 seconds.
Table 1: End-to-end runtime by phase and batch on commodity hardware (Dell G16 7630, 32 logical cores).
Phase Peano (s) Share (%) Gauss (s) Share (%)
Conjecturer 21.3 5.2 106.0 26.0
CE filter 14.6 3.6 192.5 47.3
Prover 24.6 6.0 47.1 11.6
Total (executable) 39.4 9.7 239.7 58.8
Overall wall-clock: 407 seconds
Observations
The runtime distribution shifts considerably between batches. In the Peano batch the three phases are roughly balanced; in the Gauss batch the CE filter becomes the dominant cost at 80% of the executable runtime, reflecting the increased complexity of the fact base required to screen a richer conjecture space. The prover, by contrast, remains tractable at under 50 seconds even with 64 inherited theorems broadcast into the grid. Conjecture creation accounts for the bulk of the Python-side runtime and remains the primary target for the planned migration to C++. These figures should be read as a snapshot of a initial prototype prototype; substantial headroom remains across all phases.
5 Related Work and Contributions
The field of automated reasoning has seen significant advancements from two primary directions: probabilistic models and interactive proof assistants. GL positions itself as a third paradigm, offering a deterministic and massively parallel approach to theorem discovery. This section situates GL in relation to prior work and clarifies its unique contributions.
LLMs
LLMs have demonstrated impressive capabilities in pattern matching and the reformulation of existing knowledge. They can often solve routine mathematical exercises and rephrase known proofs in a human-like manner. However, their underlying architecture is probabilistic, which presents a fundamental limitation for rigorous, multi-step deductive reasoning. LLMs do not possess an intrinsic model of logical truth; consequently, they struggle to generate novel, non-trivial proofs and are susceptible to producing plausible-sounding but logically flawed arguments ("hallucinations"). In contrast, GL is a deterministic system where every inference is a consequence of axiomatic definitions, making its outputs inherently verifiable. Recent neural approaches have started to make headway in formal theorem proving – for example, a 2025 open-source transformer model achieved about 60% success on a standard benchmark by training on millions of generated proofs [8] – but such data-intensive methods are orthogonal and complementary to GL’s logic-first architecture.
Interactive and Semi-Automated Proof Assistants.
Systems such as Lean, Rocq/Coq, and Isabelle/HOL provide expressive foundations for formalizing mathematics and verifying proofs in a small, trusted kernel [6]. Although their traditional workflow is interactive—users decompose goals, invoke tactics, and supply key intermediate lemmas—modern ecosystems expose substantial automation. Isabelle’s Sledgehammer tool, for example, exports the current goal together with a heuristic relevance slice of the local theory context to a portfolio of external automated theorem provers (ATPs) and SMT solvers (e.g., E, Vampire, Z3, cvc5); when one succeeds, Isabelle attempts proof reconstruction automatically [3]. Lean and Coq likewise support tactic automation and scripted, batch replay of large libraries. In short, contemporary ITPs can discharge many individual goals automatically once the goals and supporting libraries are in place. Coq’s ecosystem has recently incorporated “hammer” plugins (e.g. CoqHammer) that automatically attempt to prove goals by translating them to first-order logic and calling external provers [4] , often discharging nontrivial goals without manual guidance. Automated first-order provers like E have achieved remarkable success in solving individual conjectures, consistently ranking among the top systems in the annual CASC competitions[10]. Another premier first-order prover is Vampire, which has dominated many competition divisions – since 1999 it has won over 50 category titles in the “world cup” of theorem provers [7] – thanks to its highly optimized saturation engine. Satisfiability Modulo Theories (SMTs) solvers (e.g. Z3) excel at deciding logical formulas in rich theories automatically [5] , and they have become integral to many verification workflows (often being invoked behind the scenes by proof assistant tactics to discharge specific subgoals). The newest SMT solvers, such as cvc5, extend the capabilities of their predecessors (e.g. CVC4 and Z3) with support for a diverse range of theories and advanced features like higher-order reasoning and syntax-guided synthesis [2].
GL vs. ITP automation. GL differs not by rejecting such automation but by shifting the unit of work: instead of waiting for a human to pose a single conjecture, GL compiles a base set of definitions and then systematically explores their deductive closure, generating large families of conjectures and attempting proofs for all of them in parallel. Every inference step is a deterministic hash-prefix match recorded in an auditable proof graph, enabling bulk, post hoc verification outside the proving engine.
5.1 Contributions
The GL architecture introduces several novel features to the field of automated reasoning.
Definition-Initiated Bulk Theorem Generation. Where mainstream proof assistants are typically used in a goal-directed manner (a user states a conjecture and applies interactive and automated tactics—possibly calling out to external ATP/SMT tools—to discharge it), GL begins from a curated definition set. It then enumerates normalized implication templates, weaves well-typed combinations, and launches a distributed proof search over all resulting conjectures simultaneously. This enables automatic discovery of families of theorems whose statements were not explicitly posed by a user.
Deterministic and Verifiable by Construction
Every reasoning step in GL is a symbolic manipulation based on a hash-table lookup. The final output is a complete, auditable proof graph where each inference is explicitly linked to a prior definition or a previously proven theorem. This provides a level of trust and verifiability that is not available in probabilistic systems. The correctness of a proof can be independently checked by a simple, external script, without needing to trust the internal machinery of GL itself.
Massively Parallel Architecture
The core of GL is its distributed architecture, where the reasoning workload is divided among thousands or millions of independent LBs. Each block operates asynchronously, communicating only between cycles. This design allows GL to tackle computationally intensive search problems by leveraging modern parallel computing infrastructure, such as multi-core ASICs or cloud-based server clusters.
Hardware-Software Co-Design
GL is conceived as a complete system, with the algorithms and architecture designed to complement each other. This hardware-aware approach ensures that the abstract reasoning process can be efficiently mapped onto a physical or virtualized processor, making it a scalable and practical solution for large-scale mathematical exploration.
6 Future Work, Roadmap and Potential Impact
With the Gauss summation formula proven and shipped, GL has completed its first two-batch ENT run. Peano delivered algebraic manipulation; Gauss delivered logical transformation. That closes one phase. The next milestone is the Fundamental Theorem of Arithmetic (FTA).
The road to FTA is not a continuation of the same work at larger scale. It requires three infrastructure extensions and one final major algorithmic extension, arriving in a specific order dictated by their dependencies.
6.1 Pruner
GL now carries a growing mass of proven theorems. That is good for expressive power but bad for focus. The pruner’s task is to select a minimal core subset from that mass such that all remaining theorems follow directly from it. This is not merely a runtime optimization. It is a structural compression step: it removes noise from theorem selection, improves proof trace readability, and ensures that the HTML proof graph reflects the real backbone of a result rather than drowning it in nearby consequences. As the theorem base grows, this compression step becomes essential for maintaining clarity and performance.
6.2 External Proof Checker
Internal proof generation is not sufficient once GL begins carrying heavier mathematical claims. An external checker creates a hard boundary between proving and verifying. GL generates proof objects; the checker validates them independently, without access to the internal machinery of the prover. This improves trust, makes debugging cleaner, and gives the project a stronger research posture. When the prover is modified, the checker remains the stable witness. When something fails, the boundary makes it immediately clear whether the issue lies in generation or in verification.
6.3 Case Differentiation
This is the last missing tool in GL’s mathematical stack, and arguably the most consequential. Peano gave GL algebraic manipulation. Gauss gave GL logical transformation. Case differentiation is the third pillar.
A significant portion of non-trivial mathematics is not blocked by algebra or by rewrite logic — it is blocked by the need to split on a condition and drive separate proof branches under separate assumptions. GL currently follows a single linear path through a proof. Once case differentiation is in place, the system can handle proofs that depend on sign splits, divisibility splits, structural alternatives, and contradiction-driven branch closures. That is the point at which the prover stops being a corridor and becomes a machine capable of operating in branching terrain.
6.4 Incubator
The Incubator becomes possible once case differentiation is available, because it depends on that tool to generate its outputs reliably. The incubator autogenerates ground-level fact tables — the machine-generated equivalent of multiplication tables — and feeds them into the CE filter as fuel. The CE filter grows stronger in proportion to the richness of its fact stock. The incubator is therefore a throughput engine: it manufactures the raw material that allows the CE filter to operate at higher quality and higher speed, without requiring manual curation of facts as the target domain grows.
6.5 The FTA Campaign
Once the above stack is complete, the FTA campaign becomes a direct engineering effort. A route of auxiliary facts and intermediate theorems leading from the Peano axioms toward FTA will be identified and pressure-tested in sequence. If the conjecturer fails to surface a required fact, the conjecturer is extended. If the prover cannot establish it, the prover is extended. That loop repeats until the route is carried end to end.
6.6 Path to Scalability
In parallel with the mathematical extensions above, the implementation will be progressively hardened. The Python conjecturer — currently the dominant runtime cost — will be migrated to C++, recovering the multithreading headroom that Python’s GIL forecloses. The architecture will then be adapted for massively parallel cloud execution, distributing the workload across large numbers of instances on platforms such as AWS or GCP. The long-term target substrate remains a cloud of specialized ASIC chips, where the full parallelism of the Logic Block grid can be realized without constraint.
6.7 LLM Deployment
LLM Interface & Feedback Loop (post-ENT). After the ENT milestone, we will expose tool calls so an LLM can (i) submit MPL definitions and conjectures, (ii) invoke GL runs, and (iii) inspect derived theorems and proof graphs. In the reverse direction, GL becomes an LLM reasoning booster: when the model confronts a mathematical (and eventually any MPL-encodable) reasoning task, it can translate that task to MPL and ask GL to deterministically study its deductive consequences, feeding validated results back into the dialogue. GL remains an external proof engine — not a Python library import — so the LLM/GL handshake mirrors, but does not collapse into, standard Python execution flows.
7 Potential Impact (Speculative)
The architectural path sketched in this work suggests several long‑horizon applications if the performance milestones are met.
  1. 1.
    Large‑scale axiom exploration. Million‑fold throughput would permit exhaustive closure over richer theories (ENT, Algebra, selected fragments of Analysis), enabling automated conjecture mining across domains.
  2. 2.
    Deterministic back‑end for AI agents. LLMs could emit conjecture invariants or lemmas and defer all symbolic checking to GL, yielding verifiable reasoning traces.
  3. 3.
    Formal assurance at design scale. Embedding GL cores in EDA/HW verification loops could shrink overnight proof farms to on‑device checks, improving safety‑critical certification.
  4. 4.
    Cross‑theory discovery. Uniform normalisation across imported definition sets might surface unexpected links (e.g., arithmetic/algebra correspondences) that merit human follow‑up.
8 Conclusion
GL set out to answer a narrow question: whether a deterministic, definition-centric architecture can autonomously generate and verify non-trivial theorems from first principles, without human guidance at the level of individual proof steps.
The question has been answered in the affirmative, at least for the domain examined. Starting from a formalized set of Peano axioms, the system derived the foundational laws of arithmetic. Starting from those results, it derived Gauss’s summation formula. Both derivations are recorded as complete proof graphs in which every inference step is explicitly linked to its antecedent, enabling independent verification by navigating or scripting over the generated output.
Three properties of the architecture were confirmed in practice. First, every inference step is a hash-table lookup, which means every derivation is replayable and independently checkable — the proof graph is not a summary but a complete record. Second, each Logic Block holds only its local fragment of the reasoning task; blocks communicate between cycles, not during them. This clean separation is what makes the architecture parallelizable in principle and, eventually, on hardware. Third, the system was given definitions. It was not given theorems. The theorems it produced were not requested. This is the property the architecture was designed around, and it held.
The implementation is a prototype. Its performance figures are those of a prototype. The path to larger theories is identified and involves no steps that appear, at present, to be impossible.
Whether the architecture scales to the FTA and beyond remains to be established. We expect that it does.
9 List of Abbreviations
AI
Artificial Intelligence
ASIC
Application-Specific Integrated Circuit
ATP
Automated Theorem Prover
GL
Generative Logic
LB
Logic Block
RT
Run Time
HW
Hardware
ITP
Interactive Theorem Prover
LE
Logical Entity
LLM
Large Language Model
MPL
Mathematical Programming Language
SW
Software
ENT
Elementary Number Theory
FTA
Fundamental Theorem of Arithmetic
AWS
Amazon Web Service
GCP
Google Cloud Platform
CE
Counterexample
NoC
Network-on-Chip
SMT
Satisfiability Modulo Theories